# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ II.2.1 Model and eigenfunctions in altermagnetic ribbon
To illustrate the generality of our results regarding the role of boundaries in altermagnets, we start with a phenomenological model. We consider the following low-energy model [57] of a $d$-wave altermagnet, see also Fig. 1:
@@
The anisotropy of the fully spin-polarized Fermi surfaces is described by the third term in Eq. (16) with $\sigma_{z}$ being the Pauli matrix in the spin space and dimensionless parameters $|t_{1}|<1$ and $|t_{2}|<1$ describing the strength of the altermagnetism; the corresponding band structure is schematically shown in Fig. 1. The ribbon is described by the Hamiltonian (16) with $k_{x}\to-i\nabla_{x}$. We assume open boundary conditions that do not intermix spins, see Eq. (5). This and the diagonal structure of the Hamiltonian (16) allow us to treat each of the spin projections $s=\uparrow,\downarrow$ independently.
@@ II.2.2 Density of states and magnetization
where $\nu_{0}=m/(2\pi)$. Unlike antiferromagnets, the LDOS of altermagnets retains the dependence on spin, cf. Eqs. (10) and (21).
We present the spin-resolved LDOS in the middle of the ribbon, $x=L/2$, in Fig. 2. As expected from Eq. (21), there is a well-pronounced splitting of the spin-polarized DOSs in altermagnets if $t_{1}\neq 0$.
@@
$\lim_{\tilde{L}\to\infty}\!n_{s}(\tilde{x})$ $=$ $\frac{n_{0}}{\sqrt{1-t_{1}^{2}-t_{2}^{2}}}\left[1-\frac{\sqrt{1+st_{1}}}{\tilde{x}}J_{1}{\left(\frac{2\tilde{x}}{\sqrt{1+st_{1}}}\right)}\right]$
@@
$\lim_{\tilde{x}\to\infty}\nu_{s}(\tilde{\varepsilon},\tilde{x})$ $=$ $\frac{\nu_{0}}{\sqrt{1-t_{1}^{2}-t_{2}^{2}}}\left[1-\left(\frac{1+st_{1}}{\tilde{\varepsilon}+1}\right)^{1/4}\frac{1}{\sqrt{\pi}\sqrt{\tilde{x}}}\sin{\left(2\sqrt{\frac{\tilde{\varepsilon}+1}{1+st_{1}}}\tilde{x}+\frac{\pi}{4}\right)}\right]+\mathcal{O}{\left(\frac{1}{\tilde{x}}\right)},$ (26)
$\lim_{\tilde{x}\to\infty}n_{s}(\tilde{x})$ $=$ $\frac{n_{0}}{\sqrt{1-t_{1}^{2}-t_{2}^{2}}}\left[1+\frac{\left(1+st_{1}\right)^{3/4}}{\sqrt{\pi}\tilde{x}^{3/2}}\cos{\left(\frac{2}{\sqrt{1+st_{1}}}\tilde{x}+\frac{\pi}{4}\right)}\right]+\mathcal{O}{\left(\frac{1}{\tilde{x}^{2}}\right)},$ (27)
cf. Eqs. (II.1.2) and (15). While the decay with the coordinate is the same as in the antiferromagnet considered in Sec. II.1.2, the period of the oscillations is determined by the Fermi momentum, which in the present altermagnetic case reads $2k_{F}/\sqrt{1+st_{1}}$. Therefore, spin-up and spin-down charge densities are characterized by different oscillation frequencies. Hence, according to Eq. (23), an altermagnet is expected to develop oscillating magnetization.
Under the same approximations and assuming weak altermagnetic parameters, the magnetization (23) reads
$\lim_{t_{1},t_{2}\to 0}\lim_{\tilde{x}\to\infty}\!m_{z}(\tilde{x})=\frac{2t_{1}m_{0}}{\sqrt{\pi\tilde{x}}}\sin{\left(2\tilde{x}+\frac{\pi}{4}\right)}+\mathcal{O}{\left(\frac{1}{\tilde{x}^{3/2}},t_{1}^{2},t_{2}\right)},$ (28)
where $m_{0}=-g\mu_{B}n_{0}/e$. Therefore, we expect a slower $\propto 1/\sqrt{\tilde{x}}$ decay compared to the charge density oscillations. In addition, the oscillations of the local charge density and the magnetization are shifted by $\pi/2$ at small $t_{1}$.
While the LDOS and magnetization can be readily studied via local probes, it is also convenient to introduce the spatially averaged characteristics of the ribbon, which can be investigated without requiring high spatial resolution via, e.g., SQUID [51, 52, 58]. For this, we integrate the magnetization over the ribbon width:
@@
As one can clearly see from the second line of Eq. (29), the net magnetization is determined by the asymmetry between the spin-polarized Fermi surfaces in the direction perpendicular to the edge of the ribbon; the asymmetry is quantified by the eccentricity of the ellipsoidal Fermi surfaces. The same asymmetry is responsible for the different frequencies of the Friedel oscillations for spin-up and spin-down electrons.
The local $m_{z}(x)$ and net $M_{z}$ magnetizations are shown in Figs. 3(a) and 3(b), respectively. As expected, the oscillations of local magnetization contain two periods, see Eqs. (23) and (27), and decay into the bulk. For larger width of the ribbon, the oscillations in Fig. 3(b) become indiscernible and approach the result given in the last line of Eq. (29).
Figure 3: Panel (a): The spatial profile of the magnetization for a few widths of the ribbon at $t_{1}=0.5$. The periods of the oscillations are determined by two Fermi wave vectors $k_{F}/(1\pm t_{1})$. The inset shows the enlarged region near the boundary. Panel (b): The net magnetization $M_{z}$ as a function of the altermagnetic parameter $t_{1}$. The green dashed line corresponds to the approximate expression in the last line in Eq. (29). Here, $M_{0}=m_{0}/k_{F}$. In both panels, we fix $t_{2}=0$.
To summarize the results of the analytical section, we found that the spin anisotropy of the Fermi surface in altermagnets is manifested in the spin-resolved LDOS and magnetization. By using a ribbon geometry as a representative example, we found the splitting of the spin-resolved LDOS and nonvanishing magnetization for generic orientations of the altermagnetic Fermi surfaces except for the single orientation that has combined mirror and spin-flip symmetry, corresponding to a compensated interface such that the spin-polarized lobes in Fig. 1 are rotated by $\pi/4$. The splitting of the spin-resolved LDOS and resulting magnetization arise from different length scales of the Friedel oscillations quantified by $2k_{F}/\sqrt{1+st_{1}}$. For a wide ribbon, we found that the magnetization decays as $1/\sqrt{k_{F}x}$ away from the edges of the ribbon and is shifted by $\pi/2$ in phase compared to the electric charge oscillations.
@@
These analytical findings are confirmed in a lattice model for altermagnets in Sec. III. The results of the lattice model, however, are more nuanced and provide access to additional information related to the unit cell structure. In fact, we will see that the tight-binding model reveals a mechanism that induces an interface-magnetization carried by the itinerant electrons which is not captured by the analytical model, such as the interface-induced magnetization in antiferromagnets.
@@ III.1 Tight-binding $s-d$ model for antiferromagnet and altermagnet
We describe the AFM or AM ribbon on a lattice through a tight-binding model with an s-d coupling to localized anti-collinear moments. The lattice is shown in Fig. 1. This model is inspired by the microscopic model for altermagnets introduced by Ref. [59], but we consider the full four-site unit cell. This allows us to continuously introduce the lattice asymmetry which takes the AFM order ($\varepsilon_{2}=\varepsilon_{3}$) to an AM order ($\varepsilon_{2}\neq\varepsilon_{3}$). Note in particular that we take the boundary conditions in the x-direction to be open, while the y-direction is periodic.
@@
where i, j denote individual lattice sites, t describes nearest-neighbor hopping, and $J_{\text{sd}}$ is the coupling between the itinerant electrons described by $c^{\vphantom{\dagger}}/c^{{\dagger}}$ and the localized moments $\boldsymbol{S}_{i}$, the configuration of which is shown in Fig. 1. Finally, $\mu$ is the chemical potential while an additional on-site energy $\varepsilon_{\text{NM}}\in\{\varepsilon_{2},\varepsilon_{3}\}$ is present if the site is a non-magnetic site.
We now segment the lattice into four-site unit cells shown in Fig. 1 using a band basis $c_{I,\alpha,\sigma}^{\vphantom{\dagger}}/c_{I,\alpha,\sigma}^{{\dagger}}$, where I is now a unit cell index, $\alpha\in[1,2,3,4]$ denote the site index within the unit cell and where $\sigma$ is the spin-index. To take advantage of the periodicity in y, we also introduce Fourier-transformed band-operators in the y-direction,
@@
where $N_{y}^{\text{u.c.}}=N_{y}/2$ is the number of unit cells in the y direction, where $R_{y}$ is the y-coordinate of the unit cell and where the momentum $k_{y}\in(-\pi/2a,\pi/2a]$ runs over the reduced Brillouin zone corresponding to the $2\times 2$ real-space unit cell. From now on, we omit x/y subscripts on real- and momentum space indices, taking $I=I_{x}$ and $k=k_{y}$ for brevity of notation.
The total Hamiltonian in the unit cell basis is now given by
@@
where $\Delta_{x}$ is the nearest-neighbor vector ($|\Delta_{x}|=|\Delta_{y}|=2a$) between two unit cells in the x-direction, where the unit cell basis is given as
$\tilde{c}_{I,k,\sigma}=\begin{pmatrix}c_{I,k,1,\sigma}&c_{I,k,2,\sigma}&c_{I,k,3,\sigma}&c_{I,k,4,\sigma}\end{pmatrix}^{T}$ (33)
and where we have defined the coefficient matrices
@@
$h_{I,I,k,\sigma}=\begin{bmatrix}\sigma-\mu&-t&-te^{2ika}-t&0\\ -t&\varepsilon_{2}-\mu&0&-te^{2ika}-t\\ -t\mathrm{e}^{-2ika}-t&0&\varepsilon_{3}-\mu&-t\\ 0&-te^{-2ika}-t&-t&-\sigma-\mu\end{bmatrix},\qquad h_{I,I+\Delta_{x},k,\sigma}=-t\begin{bmatrix}0&\delta_{\Delta_{x},-\hat{x}}&0&0\\ \delta_{\Delta_{x},\hat{x}}&0&0&0\\ 0&0&0&\delta_{\Delta_{x},-\hat{x}}\\ 0&0&\delta_{\Delta_{x},\hat{x}}&0\end{bmatrix}$ (34)
@@
Figure 4: The spectral function $A_{\sigma}(k_{x},E)$ is shown for $k_{y}=0$ in the model with translational invariance in both x and y. The NM band structure is shown in (a) where the energy states between $E\sim-2t$ and $E\sim 2t$ are backfolded into the reduced Brillouin zone (RBZ) due to the unit cell segmentation. Introducing a non-zero $J_{\text{sd}}=t$, gaps open at the AFM zone edge (b). If one additionally breaks the equivalence between the two non-magnetic sites, a spin-split altermagnetic band structure emerges (c), in this case by setting $\varepsilon_{3}=100t$. The insets show the Fermi surface at the chemical potential denoted by the dotted line.
@@
We have explicitly set $\boldsymbol{S}_{i}\propto\hat{\boldsymbol{z}}$. Moreover, we note that the antiferromagnetic state can be modelled with a simpler lattice, namely a 2-site unit cell where the normal atoms in our model have been omitted. We have confirmed that the essential physical mechanisms underlying the interface-induced magnetization, which will be explained below, are also present in such a model. To elucidate the spectral information of this model, the spectral functions are plotted for a bulk model (translational invariant in both x and y) for $k_{x}$ vs $E$ in Fig. 4 showing the effect of a non-zero $J_{\text{sd}}$ and a broken equivalence between $\varepsilon_{2}$ and $\varepsilon_{3}$ on the spectral properties of the model.
@@
Observables such as magnetization and charge density are now readily obtainable after diagonalization of Eq. (32). Depending on the interplay between spin-dependent ($J_{\text{sd}}$) and spin-independent ($\varepsilon_{2}$, $\varepsilon_{3}$) on-site couplings within the unit cell, the effect of the interfaces on the system density can vary significantly. In the following, we begin by discussing how the introduction of non-magnetic on-site potentials alters the system eigenfunctions and the subsequent consequences on the system density. We then move on to discuss how these effects give rise to interface-induced magnetism in antiferromagnets and altermagnets.
@@
III.2 Wave function shift, magnitudes and Friedel oscillations: illustration of concepts using a normal metal
To provide the basis of our discussion of the magnetization and charge oscillations in ribbons of antiferromagnets and altermagnets, let us start with clarifying the nuances of the lattice treatment of a regular metal.
@@
The characterization of magnetization and edge effects revolves around the study of the spin-resolved electron density and how it varies across the lattice. The eigenfunctions of our lattice, $\psi_{n,k,\sigma}(r_{I},\alpha)$, are obtained from the column vectors of the matrix $U$ diagonalizing the Hamiltonian, and are resolved at the unit cell level. Each eigenfunction is described by three indices $k,n,\sigma$ and is a function of the x-coordinate I and the site index $\alpha$ within each unit cell. The spin- and $\alpha$-dependent density at x-position $r_{I}$ in the ribbon is defined as
@@
Figure 5: NM densities obtained by a sum over the two lowest energy harmonics for $k=0$ shown in the unit cell-resolved (a) and row-resolved (b) case. Analogous plots in (c) and (d) are shown in the case of a non-zero on-site coupling $\varepsilon_{2}=\delta$, $\varepsilon_{3}=-\delta$ for $\delta\ll t$. (a) In the NM, the densities are pair-wise equivalent with the ones corresponding to sites in the left column (blue) and right column (red) equal. The left and right column eigenfunctions are shifted by the lattice constant $a$, corresponding to their relative displacement within the unit cell. (b) The row-resolved density can be illustrated by adding the densities for $\alpha=1$ and $\alpha=2$ for the row 1 and $\alpha=3$ and $\alpha=4$ for the 2. row. (c) A change in the on-site potentials changes the magnitude of the eigenfunction entries corresponding to the particular site, lifting the previous equivalence between the densities at different sites $\alpha$. The probability amplitude at $\alpha=2$ sites is suppressed due to $\varepsilon_{2}=\delta$, while the negative potential $\varepsilon_{3}=-\delta$ enhances the probability amplitude at the $\alpha=3$ sites. The row-resolved densities are given by the densities of unit cell sites within the same row, shifted relatively by a lattice constant $a$. Due to the induced magnitude difference, the system density becomes a $\pi$-periodic density wave.
@@
Before moving on, we briefly mention that in the following, we will use the terminology unit cell-resolved and row-resolved extensively. Row-resolved refers to quantities that are functions of individual lattice site indices $(i,k)$ across one of the two distinct rows on our ribbon (first and second row in the unit cell, see Fig. 1), while unit cell-resolved refers to quantities which are functions of unit cell indices $(I,\alpha,k)$.

### v3

@@ II.2.1 Model and eigenfunctions in altermagnetic ribbon
To illustrate the generality of our results regarding the role of boundaries in altermagnets, we start with continuum model obtained from symmetry considerations [57]; such a model can also be derived from the lattice model defined in Sec. III.1, see Appendix A for the details. We use the following Hamiltonian describing 2D $d$-wave altermagnets:
@@
The anisotropy of the fully spin-polarized Fermi surfaces is described by the third term in Eq. (15) with $\sigma_{z}$ being the Pauli matrix in the spin space and dimensionless parameters $|t_{1}|<1$ and $|t_{2}|<1$ describing the strength of the altermagnetism; the corresponding band structure is schematically shown by the blue and red ellipses in Fig. 1(a) at $t_{1}\neq 0$ and $t_{2}=0$. The case with $t_{1}=0$ and $t_{2}\neq 0$ would correspond to a rotation of both the blue and red ellipses by $\pi/4$. In the latter case, the interface does not break the symmetry between the spin-polarized Fermi surfaces. The ribbon is described by the Hamiltonian (15) with $k_{x}\to-i\nabla_{x}$. We assume open boundary conditions that do not intermix spins, see Eq. (4). This and the diagonal structure of the Hamiltonian (15) allow us to treat each of the spin projections $s=\uparrow,\downarrow$ independently.
@@ II.2.2 Density of states and magnetization
where $\nu_{0}=m/(2\pi)$ is the bulk DOS in 2D. Unlike antiferromagnets, the LDOS of altermagnets retains the dependence on spin, cf. Eqs. (9) and (20).
We present the spin-resolved LDOS in the middle of the ribbon, $x=L/2$, in Fig. 2. As expected from Eq. (20), there is a well-pronounced splitting of the spin-polarized DOSs in altermagnets if $t_{1}\neq 0$, i.e., when the altermagnetic lobes are perpendicular to the interface, see Fig. 1. No splitting is observed for $t_{1}=0$ and $t_{2}\neq 0$ when the symmetry between the spin-polarized Fermi surfaces is preserved by the interface.
@@
$\lim_{\tilde{L}\to\infty}\!n_{s}(\tilde{x})$ $=$ $\frac{n_{0}}{\sqrt{1-t_{1}^{2}-t_{2}^{2}}}\left[1-\frac{\sqrt{1+st_{1}}}{\tilde{x}}J_{1}{\left(\frac{2\tilde{x}}{\sqrt{1+st_{1}}}\right)}\right],$
@@
$\nu_{s}(\tilde{\varepsilon},\tilde{x})$ $=$ $\frac{\nu_{0}}{\sqrt{1-t_{1}^{2}-t_{2}^{2}}}\left[1-\left(\frac{1+st_{1}}{\tilde{\varepsilon}+1}\right)^{1/4}\frac{1}{\sqrt{\pi\tilde{x}}}\sin{\left(2\sqrt{\frac{\tilde{\varepsilon}+1}{1+st_{1}}}\tilde{x}+\frac{\pi}{4}\right)}\right]+\mathcal{O}{\left(\frac{1}{\tilde{x}}\right)},$ (25)
$n_{s}(\tilde{x})$ $=$ $\frac{n_{0}}{\sqrt{1-t_{1}^{2}-t_{2}^{2}}}\left[1+\frac{\left(1+st_{1}\right)^{3/4}}{\sqrt{\pi}\tilde{x}^{3/2}}\cos{\left(\frac{2}{\sqrt{1+st_{1}}}\tilde{x}+\frac{\pi}{4}\right)}\right]+\mathcal{O}{\left(\frac{1}{\tilde{x}^{2}}\right)},$ (26)
cf. Eqs. (II.1.2) and (II.1.2). While the decay with the coordinate is the same as in the antiferromagnet considered in Sec. II.1.2, the period of the oscillations is determined by the Fermi momentum, which in the present altermagnetic case reads as $2k_{F}/\sqrt{1+st_{1}}$. Therefore, spin-up and spin-down charge densities are characterized by different oscillation frequencies. Hence, according to Eq. (22), an altermagnet is expected to develop oscillating magnetization.
Under the same approximations and assuming weak altermagnetic parameters, $t_{1}\ll 1$ and $t_{2}\ll 1$, the magnetization (22) reads
$m_{z}(\tilde{x})\approx\frac{2t_{1}m_{0}}{\sqrt{\pi\tilde{x}}}\sin{\left(2\tilde{x}+\frac{\pi}{4}\right)},$ (27)
where $m_{0}=-g\mu_{B}n_{0}/e$. In writing Eq. (27), we expanded in $1/\tilde{x}$, $t_{1}$, and $t_{2}$ and kept only the leading-order nontrivial term. Therefore, we expect a slower $\propto 1/\sqrt{\tilde{x}}$ decay compared to the charge density oscillations. In addition, the oscillations of the local charge density and the magnetization are shifted by $\pi/2$ at small $t_{1}$.
While the LDOS and magnetization can be readily studied via local probes, it is also convenient to introduce the spatially averaged characteristics of the ribbon, which can be investigated without requiring high spatial resolution via, e.g., SQUID [51, 58, 59, 52, 60]. For this, we integrate the magnetization over the ribbon width:
@@
As one can see from the second line of Eq. (28), the net magnetization is determined by the asymmetry between the spin-polarized Fermi surfaces in the direction perpendicular to the edge of the ribbon; the asymmetry is quantified by the eccentricity of the ellipsoidal Fermi surfaces. The same asymmetry is responsible for the different frequencies of the Friedel oscillations for spin-up and spin-down electrons.
The local $m_{z}(x)$ and net $M_{z}$ magnetizations are shown in Figs. 3(a) and 3(b), respectively. As expected, the oscillations of local magnetization contain two periods, see Eqs. (22) and (26), and decay into the bulk. For a larger width of the ribbon, the oscillations in Fig. 3(b) become indiscernible and approach the result given in the last line of Eq. (28).
Figure 3: (a) The spatial profile of the magnetization for a few widths of the ribbon at $t_{1}=0.5$. The periods of the oscillations are determined by two Fermi wave vectors $k_{F}/(1\pm t_{1})$. The inset shows the enlarged region near the boundary. Panel (b) The magnetization $M_{z}$ obtained by integrating over the ribbon width as a function of the altermagnetic parameter $t_{1}$. The green dashed line corresponds to the approximate expression in the last line in Eq. (28). Here, $M_{0}=m_{0}/k_{F}$. In both panels, we fix $t_{2}=0$.
To provide numerical estimates of the induced magnetization, we use the gyromagnetic ratio $g=2$ and estimate the effective mass from the tight-binding model expansion as $m=1/(2a^{2}t)$ (see, e.g., Eq. (39)) with $a=5~{}\AA$ being the typical lattice constant and $t=1~{}\mbox{eV}$ being the hopping parameter. The characteristic magnetization $m_{0}$ and corresponding magnetic polarization $\mathcal{P}_{m_{0}}$ are
$m_{0}\approx 6\times 10^{-3}\mu_{B}\,\frac{\mu}{1~{}\mbox{eV}},\quad\mathcal{P}_{m_{0}}\approx 0.1\,\frac{\mu}{1~{}\mbox{eV}}~{}\mu\mbox{G}.$ (29)
The characteristic magnitude of the integrated over the ribbon width magnetization (i.e., the flux per unit length) $M_{0}=m_{0}/k_{F}$, see Eq. (28), and the corresponding integrated magnetic polarization are estimated as
$M_{0}\approx 3\times 10^{-2}\,t_{1}\mu_{B}\,\frac{\mu}{1~{}\mbox{eV}}\,\AA,\quad\mathcal{P}_{M_{0}}\approx 0.4t_{1}\,\frac{\mu}{1~{}\mbox{eV}}~{}\mu\mbox{G}\,\AA.$ (30)
[…]
To summarize the results of this section, we found that the spin anisotropy of the Fermi surface in altermagnets is manifested in the spin-resolved LDOS and magnetization. By using a ribbon geometry as a representative example, we found the splitting of the spin-resolved LDOS and nonvanishing magnetization for generic orientations of the altermagnetic Fermi surfaces except for the single orientation that has combined mirror and spin-flip symmetry, corresponding to a compensated interface such that the spin-polarized lobes in Fig. 1(a) are rotated by $\pi/4$. The splitting of the spin-resolved LDOS and resulting magnetization arise from different length scales of the Friedel oscillations quantified by $2k_{F}/\sqrt{1+st_{1}}$. For a wide ribbon, we found that the magnetization decays as $1/\sqrt{k_{F}x}$ away from the edges of the ribbon and is shifted by $\pi/2$ in phase compared to the electric charge oscillations.
@@
These analytical findings are confirmed in a lattice model for altermagnets in Sec. III. The results of the lattice model, however, are more nuanced and provide access to additional information related to the unit-cell structure. In fact, we will see that the tight-binding model reveals a mechanism that induces an interface magnetization carried by the itinerant electrons which is not captured by the analytical model, such as the interface-induced magnetization in antiferromagnets.
@@ III.1 Tight-binding s-d model for antiferromagnet and altermagnet
We describe the AFM or AM ribbon on a lattice through a tight-binding model with an s-d coupling to localized anti-collinear moments. The lattice is shown in Fig. 1(b). This model is inspired by the microscopic model for altermagnets introduced by Ref. [61], but we consider the full four-site unit-cell. This allows us to continuously introduce the lattice asymmetry which takes the AFM order ($\varepsilon_{2}=\varepsilon_{3}$) to the AM order ($\varepsilon_{2}\neq\varepsilon_{3}$). Note in particular that we take the boundary conditions in the x-direction to be open, while the y-direction is periodic.
@@
where i, j denote individual lattice sites, t describes nearest-neighbor hopping, and $J_{\text{sd}}$ is the coupling between the itinerant electrons described by $c^{\vphantom{\dagger}}/c^{{\dagger}}$ and the localized moments $\boldsymbol{S}_{i}$, the configuration of which is shown in Fig. 1(b). Finally, $\mu$ is the chemical potential while an additional on-site energy $\varepsilon_{\text{NM}}\in\{\varepsilon_{2},\varepsilon_{3}\}$ is present if the site is nonmagnetic.
We now segment the lattice into four-site unit-cells shown in Fig. 1(b) using a band basis $c_{I,\alpha,\sigma}^{\vphantom{\dagger}}/c_{I,\alpha,\sigma}^{{\dagger}}$, where I is now a unit-cell index, $\alpha\in[1,2,3,4]$ denote the site index within the unit-cell and where $\sigma$ is the spin-index. To take advantage of the periodicity in y, we also introduce Fourier-transformed band-operators in the y-direction,
@@
where $N_{y}^{\text{u.c.}}=N_{y}/2$ is the number of unit-cells in the y direction, where $R_{y}$ is the y-coordinate of the unit-cell and where the momentum $k_{y}\in(-\pi/2a,\pi/2a]$ runs over the reduced Brillouin zone corresponding to the $2\times 2$ real-space unit-cell and $a$ is the lattice constant. From now on, we omit x/y subscripts on real- and momentum space indices, taking $I=I_{x}$ and $k=k_{y}$ for brevity of notation.
The total Hamiltonian in the unit-cell basis is now given by
@@
where $\Delta_{x}$ is the nearest-neighbor vector ($|\Delta_{x}|=|\Delta_{y}|=2a$) between two unit-cells in the x-direction, the unit-cell basis is given as
$\tilde{c}_{I,k,\sigma}=\begin{pmatrix}c_{I,k,1,\sigma}&c_{I,k,2,\sigma}&c_{I,k,3,\sigma}&c_{I,k,4,\sigma}\end{pmatrix}^{T},$ (34)
and where we defined the coefficient matrices
@@
$h_{I,I,k,\sigma}=\begin{bmatrix}-J_{\text{sd}}\sigma-\mu&-t&-te^{2ika}-t&0\\ -t&\varepsilon_{2}-\mu&0&-te^{2ika}-t\\ -t\mathrm{e}^{-2ika}-t&0&\varepsilon_{3}-\mu&-t\\ 0&-te^{-2ika}-t&-t&J_{\text{sd}}\sigma-\mu\end{bmatrix},\qquad h_{I,I+\Delta_{x},k,\sigma}=-t\begin{bmatrix}0&\delta_{\Delta_{x},-\hat{x}}&0&0\\ \delta_{\Delta_{x},\hat{x}}&0&0&0\\ 0&0&0&\delta_{\Delta_{x},-\hat{x}}\\ 0&0&\delta_{\Delta_{x},\hat{x}}&0\end{bmatrix}.$ (35)
@@
Figure 4: The spectral function $A_{\sigma}(k_{x},E)$ (see Appendix B for definition) is shown for $k_{y}=0$ in the lattice model with translational invariance in both x and y. The normal metal (NM) band structure is shown in (a) where the energy states between $E\sim-2t$ and $E\sim 2t$ are backfolded into the reduced Brillouin zone (RBZ) due to the unit-cell segmentation. Introducing a non-zero $J_{\text{sd}}=t$, gaps open at the AFM zone edge (b). If one additionally breaks the equivalence between the two nonmagnetic sites, a spin-split altermagnetic band structure emerges (c). By setting $\varepsilon_{3}$ to a large value relative to other magnitudes in the system, $\varepsilon_{3}=100t$, we make this site effectively unavailable. The insets show the Fermi surface at the chemical potential denoted by the dotted line.
@@
We explicitly set $\boldsymbol{S}_{i}\propto\hat{\boldsymbol{z}}$. Moreover, we note that the antiferromagnetic state can be modeled with a simpler lattice, namely a two-site unit-cell that contains only magnetic atoms. We confirmed that the essential physical mechanisms underlying the interface-induced magnetization, which will be explained below, are also present in such a model.
To elucidate the spectral information of this model, the spectral functions (see Appendix B for definition) are plotted for the bulk model (translational invariant in both x and y) as the function of the momentum $k_{x}$ and the energy $E$ in Fig. 4. The latter shows the effects of a non-zero $J_{\text{sd}}$ and a broken equivalence between $\varepsilon_{2}$ and $\varepsilon_{3}$ on the spectral properties of the model.
@@
Observables such as magnetization and charge density are now readily obtainable after diagonalization of Eq. (33). Depending on the interplay between spin-dependent ($J_{\text{sd}}$) and spin-independent ($\varepsilon_{2}$, $\varepsilon_{3}$) on-site couplings within the unit-cell, the effect of the interfaces on the system density can vary significantly. In the following, we begin by discussing how the introduction of nonmagnetic on-site potentials alters the system eigenfunctions and the subsequent consequences on the system density. Then, we discuss how these effects give rise to interface-induced magnetism in antiferromagnets and altermagnets.
@@
III.2 Wave function shift, magnitudes, and Friedel oscillations: illustration of concepts using a normal metal
To provide the basis of our discussion of the magnetization and charge oscillations in ribbons of antiferromagnets and altermagnets, let us start by clarifying the nuances of the lattice treatment of a regular metal. The purpose of this section is to identify the key mechanisms at play in the microscopic model concerning spin-independent density modulations. With an understanding of the mechanisms that determine the oscillation pattern of the electron density near an interface, we can proceed to study the consequences of spin-dependent potentials for antiferromagnets and altermagnets.
@@
The characterization of magnetization and edge effects revolves around the study of the spin-resolved electron density and how it varies across the lattice. The eigenfunctions of our lattice, $\psi_{n,k,\sigma}(r_{I},\alpha)$, are obtained from the column vectors of the matrix $U$ diagonalizing the Hamiltonian, and are resolved at the unit-cell level. Each eigenfunction is described by three indices $k,n,\sigma$ and is a function of the x-coordinate I and the site index $\alpha$ within each unit-cell. The spin- and $\alpha$-dependent density at x-position $r_{I}$ in the ribbon is defined as
@@
Figure 5: The NM densities obtained by summing over the two lowest energy harmonics for $k=0$ shown in the unit-cell-resolved (a) and row-resolved (b) case. Analogous plots in (c) and (d) are shown in the case of a non-zero on-site coupling $\varepsilon_{2}=\delta$, $\varepsilon_{3}=-\delta$ for $\delta\ll t$. (a) In the NM, the densities are pair-wise equivalent with the ones corresponding to sites in the left column (blue) and right column (red) equal. The left and right column eigenfunctions are shifted by the lattice constant $a$, corresponding to their relative displacement within the unit-cell. (b) The row-resolved density can be illustrated by adding the densities for $\alpha=1$ and $\alpha=2$ for the row 1 and $\alpha=3$ and $\alpha=4$ for the row 2. (c) A change in the on-site potentials changes the magnitude of the eigenfunction entries corresponding to the particular site, lifting the previous equivalence between the densities at different sites $\alpha$. The probability amplitude at $\alpha=2$ sites is suppressed due to $\varepsilon_{2}=\delta$, while the negative potential $\varepsilon_{3}=-\delta$ enhances the probability amplitude at the $\alpha=3$ sites. The row-resolved densities are given by the densities of unit-cell sites within the same row, shifted relatively by a lattice constant $a$. Due to the induced magnitude difference, the system density becomes a $\pi$-periodic density wave.
@@
Before moving on, we briefly mention that in the following, we will use the terminology unit-cell-resolved and row-resolved extensively. Row-resolved refers to quantities that are functions of individual lattice-site indices $(i,k)$ across one of the two distinct rows on our ribbon [first and second rows in the unit-cell, see Fig. 1(b)], while unit-cell-resolved refers to quantities which are functions of unit-cell indices $(I,\alpha,k)$.